OpenStudy (anonymous):

solve 2x^2-8x=-7 (quadratic equation)

OpenStudy (anonymous):

@Firejay5 @kewlgeek555

OpenStudy (firejay5):

+ 7 to both side to make the right side equal to 0

OpenStudy (firejay5):

Divide by 2 ----> x^2 - 4x + 7/2 = 0, then subtract 7/2 from both sides to get x^2 - 4x = 7/2, then add 4 to both sides to get x^2 = - 1/2

OpenStudy (firejay5):

factor out the left size (x - 2)^2 = 1/2, take the square root of both sides to get x -2 = 1/sqrt of 2 or x - 2 = -1/sqrt of 2 <---- solve both equations and your answer is x = 2+/- 1/sqrt of 2

OpenStudy (anonymous):

your the best(:

OpenStudy (firejay5):

no problem, go to Wolframalpha.com/; it'll help you. http://www.wolframalpha.com/input/?i=2x%5E2+-+8x+%2B+7+%3D+0&lk=4&num=1

OpenStudy (kewlgeek555):

I think mathway.com is also helpful @savannaxx_